# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ II-B Convergence Analysis: Non-convex
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},$ (34)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq$ $\frac{4(f(\bar{{\bf{x}}}^{0})-f^{\star})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\big{(}\chi^{3}+256\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},$ (35)
@@
Suppose that Assumptions 1, 2, and 4 hold. If $\beta=1$, $\chi$ satisfies that $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{(1-\lambda_{2})T}\right).$ (36)
Proof See Appendix VII. $\Box$ When $p<\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$. Choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$. Since in each iteration we trigger communication with probability $p$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ is bounded by:
$p<\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{1-\lambda_{2}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{1}{\epsilon}\right).$
@@
$p\geq\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{1}{\epsilon}\right).$
II-C Convergence Analysis: Convex
@@
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\},$ (37)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{L\alpha^{2}\sigma^{2}\big{(}\chi^{3}+384\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}}.$ (38)
Proof See Appendix VIII. $\Box$
@@
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu\geq 0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{(1-\lambda_{2})T}\right).$ (39)
Proof See Appendix IX. $\Box$ Similar as the analysis of non-convex setting, when $p<\lambda_{2}$ and choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}\leq\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$, for any desired accuracy $\epsilon>0$, the expected communication complexity is bounded by
$p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{1-\lambda_{2}}\frac{\sigma\sqrt{L}}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{L}{\epsilon}\right).$
@@
$p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma\sqrt{L}}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{L}{\epsilon}\right).$
II-D Convergence Analysis: Strongly Convex
@@
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\frac{\sqrt[3]{4\mu(1-\gamma)}}{L}\right\},$ (40)
where $\gamma=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq$ $\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{t}\Big{(}\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}\Big{)}+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{12\mu(1-\lambda_{2})\chi}.$ (41)
Proof See Appendix X. $\Box$
@@
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu>0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{\mu(1-\lambda_{2})}{\chi L^{2}}\right)$ such that
$\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{\sigma^{2}\chi}{(1-\lambda_{2})T^{2}}+\mathrm{exp}\Big{[}-\frac{(1-\lambda_{2})T}{\chi}\Big{]}\right).$ (42)
Proof See Appendix XI. $\Box$
Similar as the analysis of non-convex and convex settings, we have $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}\leq\mathcal{O}(\nicefrac{{1}}{{(1-\lambda_{2})}})$ if $p<\lambda_{2}$ and $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}=\mathcal{O}(1)$ if $p\geq\lambda_{2}$. Thus, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$ is bounded by
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p}{{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{(1-\lambda_{2})^{2}}\right),\ p\in(0,\lambda_{2}),$
@@
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{1-\lambda_{2}}\right),\ p\in[\lambda_{2},1].$
III Proof Theorem 2

### v4

@@ VII-B Convergence Analysis: Non-convex
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},$ (35)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]\leq$ $\frac{4(f(\bar{{\bf{x}}}^{0})-f^{\star})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\big(\chi^{3}+256\chi(2\chi^{2}+p(1-p))\big)}{2p(1-\lambda_{2})\chi^{2}},$ (36)
@@
Suppose that Assumptions 1, 2, and 4 hold. If $\beta=1$, $\chi$ satisfies that $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha\leq\mathcal{O}\left(\frac{p(1-\lambda_{2})}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right).$ (37)
Proof See Appendix XII. $\Box$
Let $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$. Solving
$\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right)<\epsilon,$
we have the iteration complexity of Algorithm 1 for the non-convex setting is
$T=\mathcal{O}\left(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\chi}}{\sqrt{p(1-\lambda_{2})}}\frac{\sigma}{\epsilon^{\frac{3}{2}}}+\frac{\chi}{p(1-\lambda_{2})\epsilon}\right).$
Since in each iteration we trigger communication with probability $p$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ is bounded by:
$pT=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p\chi}}{\sqrt{(1-\lambda_{2})}}\frac{\sigma}{\epsilon^{\frac{3}{2}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right),$
i.e., the expected communication complexity (10) holds.
When $p<\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$. Choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ is bounded by:
$p<\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{1-\lambda_{2}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{1}{\epsilon}\right).$
@@
$p\geq\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{1}{\epsilon}\right).$
VII-C Convergence Analysis: Convex
@@
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\},$ (38)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{L\alpha^{2}\sigma^{2}\big(\chi^{3}+384\chi(2\chi^{2}+p(1-p))\big)}{2p(1-\lambda_{2})\chi^{2}}.$ (39)
Proof See Appendix XIII. $\Box$
@@
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu\geq 0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha\leq\mathcal{O}\left(\frac{p(1-\lambda_{2})}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right).$ (40)
Proof See Appendix XIV. $\Box$
Solving
$\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right)<\epsilon,$
we have the iteration complexity of Algorithm 1 for the convex setting is
$T=\mathcal{O}\left(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\chi}}{\sqrt{p(1-\lambda_{2})}}\frac{\sigma}{\epsilon^{\frac{3}{2}}}+\frac{\chi}{p(1-\lambda_{2})\epsilon}\right).$
Since in each iteration we trigger communication with probability $p$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ is bounded by:
$pT=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p\chi}}{\sqrt{(1-\lambda_{2})}}\frac{\sigma}{\epsilon^{\frac{3}{2}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right),$
i.e., the expected communication complexity (10) holds.
When $p<\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$. Choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]$ is bounded by:
$p<\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{1-\lambda_{2}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{1}{\epsilon}\right).$
@@
$p\geq\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{1}{\epsilon}\right).$
VII-D Convergence Analysis: Strongly Convex
@@
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\frac{\sqrt[3]{4\mu(1-\gamma)}}{L}\right\},$ (41)
where $\gamma=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}<1$, it holds that
$\mathbb{E}\!\left[\big\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big\|^{2}\right]\leq$ $\Big(1-\frac{\alpha\mu}{4}\Big)^{t}\Big(\big\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big\|^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{p(1-\lambda_{2})}\Big)+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{12\mu p(1-\lambda_{2})\chi}.$ (42)
Proof See Appendix XV. $\Box$
@@
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu>0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha\leq\mathcal{O}\left(\frac{\mu p(1-\lambda_{2})}{\chi L^{2}}\right)$ such that
$\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{\sigma^{2}\chi}{p(1-\lambda_{2})T^{2}}+\mathrm{exp}\Big[-\frac{p(1-\lambda_{2})T}{\chi}\Big]\right).$ (43)
Proof See Appendix XVI. $\Box$
Similar as the analysis of non-convex and convex settings, we have $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}=\mathcal{O}(\nicefrac{{1}}{{(1-\lambda_{2})}})$ if $p<\lambda_{2}$ and $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}=\mathcal{O}(1)$ if $p\geq\lambda_{2}$. Thus, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$ is bounded by
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{\sqrt{p}}{{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{(1-\lambda_{2})^{2}}\right),\ p\in(0,\lambda_{2}),$
@@
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{\sqrt{p}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{1-\lambda_{2}}\right),\ p\in[\lambda_{2},1].$
VIII Proof Theorem 2